# Superconductivity from spin-canting fluctuations in rhombohedral graphene

### v2

@@ E.3 Effect of gapping out the magnons with a Zeeman field
$g^{(2)}_{\rm l,x}(\nu,{\boldsymbol{q}}=0)\sim\frac{(V+J)\cos^{2}\theta_{0}}{32z_{s}D}\left(\frac{1}{\pi\sqrt{a|B_{\parallel}|}}-\frac{\nu^{2}}{12\pi(a|B_{\parallel}|)^{3/2}}\right).$ (91)
To capture the behavior above we use a modified low-energy and momentum form of the total pairing interaction [compare to Eq. (48)],
$g_{\Phi}^{(2)}(q)=g_{\rm l}^{(2)}(q)+g_{\rm x}^{(2)}(q)\sim\frac{I}{\sqrt{\nu^{2}+c_{s}^{2}{\boldsymbol{q}}^{2}+\pi^{2}a|B_{\parallel}|}},$ (92)
which reproduces the leading low-frequency behavior above for ${\boldsymbol{q}}=0$ (up to a prefactor of order 1 for the $\sim\nu^{2}$ term). Integrating over the momentum parallel to the Fermi surface leads to the averaged interaction
[…]
The logarithmic divergence in the averaged pairing interaction is similarly cut-off at low frequencies, compared to Eq. (50), due to the magnon gap $\Delta_{\rm mag}$.

### v3

@@ Appendix D Derivation of magnon-mediated interaction
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $\frac{\tilde{J}\cos^{2}\theta_{0}}{\left[2\tilde{J}\kappa\cos^{2}\theta_{0}+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa\cos^{2}\theta_{0}D{{\boldsymbol{q}}}^{2}}$
$-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{\tilde{J}\cos^{2}\theta_{0}}{4\tilde{J}^{2}\sin^{2}\theta_{0}+\left[2\tilde{J}\kappa\left(1+\sin^{2}\theta_{0}\right)+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa(1+\sin^{2}\theta_{0})D{{\boldsymbol{q}}}^{2}},$ (77)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $-\frac{\tilde{J}\cos^{2}\theta_{0}+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}+i\gamma\nu}{\left[2\tilde{J}\kappa\cos^{2}\theta_{0}+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa\cos^{2}\theta_{0}D{{\boldsymbol{q}}}^{2}}$
$-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{\tilde{J}(1+\sin^{2}\theta_{0})+\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}+i\gamma\nu}{4\tilde{J}^{2}\sin^{2}\theta_{0}+\left[2\tilde{J}\kappa\left(1+\sin^{2}\theta_{0}\right)+\gamma^{2}\right]\nu_{n}^{2}+2\tilde{J}\kappa(1+\sin^{2}\theta_{0})D{{\boldsymbol{q}}}^{2}},$ (78)
For generic canting $20<\theta_{0}<\pi/$ 2 we have
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $\frac{\cos^{2}\theta_{0}}{2z_{s}\kappa}\frac{1}{\nu^{2}+c_{s}^{2}{{\boldsymbol{q}}}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{\cos^{2}\theta_{0}}{2z_{a}\kappa}\frac{1}{\Omega_{a-}^{2}+\nu^{2}+c_{a}^{2}{{\boldsymbol{q}}}^{2}},$ (79)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $\frac{\cos^{2}\theta_{0}}{2z_{s}\kappa}\frac{1}{\nu^{2}+c_{s}^{2}{{\boldsymbol{q}}}^{2}}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1+\sin^{2}\theta_{0}}{2z_{a}\kappa}\frac{1}{\Omega_{a-}^{2}+\nu^{2}+c_{a}^{2}{{\boldsymbol{q}}}^{2}},$ (80)
Here, we have defined
$c_{s}$ $=\left(\frac{D}{\kappa z_{s}}\right)^{1/2}\cos\theta_{0},\quad z_{s}=\cos^{2}\theta_{0}+\frac{V-J}{2J},$ (81)
$c_{a}$ $=\left(\frac{D}{\kappa z_{a}}\right)^{1/2}(1+\sin^{2}\theta_{0})^{1/2},\quad z_{a}=(1+\sin^{2}\theta_{0})+\frac{V-J}{2J}$ (82)
$\Omega_{a-}$ $=\sqrt{\frac{2\tilde{J}}{z_{a}\kappa}}\sin\theta_{0}\sim\left(\frac{J\sin\theta_{0}}{V}\right)\sqrt{\frac{1}{V\kappa}}\sim\left(\frac{J\sin\theta_{0}}{V}\right)\epsilon_{F},\quad\text{when $V\gg J$}.$ (83)
Here we have used the identity $\gamma^{2}/\tilde{J}\kappa=(V-J)/J$, which follows from Eq. (51), and assumed that $\bar{m}\sim\epsilon_{F}$ as in the main text.
For the case of $\theta_{0}=\pi/2$ which is achieved in the absence of Ising SOC, we have
$g^{+-,+-}_{\xi_{1},\xi_{2}}(q)=$ 00 0 (84)
$g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim$ $-\frac{1}{\kappa\nu^{2}+D{{\boldsymbol{q}}}^{2}-i\gamma\nu}-\tau_{\xi_{1}}\tau_{\xi_{2}}\frac{1}{2z_{a}\kappa}\frac{1}{\Omega_{a-}^{2}+\nu^{2}+c_{a}^{2}{{\boldsymbol{q}}}^{2}}.$ (85)
Here, in $g^{-+,+-}$, the first term describes the contribution of symmetric magnon modes, which merge into one branch of quadratic magnon at $\theta_{0}=0$. The second term describes the contribution of the gapped antisymmetric modes. At small $q$ and $\nu$, Eq. (85) reduces to $g^{-+,+-}_{\xi_{1},\xi_{2}}(q)\sim-\frac{1}{-i\gamma\nu+D{\boldsymbol{q}}^{2}}$.